The supply and demand functions for a product are π‘žπ‘  = 𝑝 2 βˆ’ 300 π‘Žπ‘›π‘‘ π‘žπ‘‘ = 𝑝 2 βˆ’ 30𝑝 + 2200 . Determine the market equilibrium price and quantity.



1
Expert's answer
2021-05-26T14:17:27-0400

In equilibrium "q_s=q_d"


"p^2-300=p^2-30p+2200"

"30p=2500"

"p=\dfrac{250}{3}\approx83.33"

"q_s=(\dfrac{250}{3})^2-300\approx6644"

"q_d\approx6644"